OpenStudy (anonymous):

A boat is sailing due north, as indicated by its compass needle, in a location where Earth's magnetic field is 2.0 X 10-5 T. The boat’s captain inadvertently places his radio on the shelf directly above the compass. If the 5.0 A current carrying wire of the radio is aligned in a north-south direction and is 30 cm directly above the compass, how much from due north with the compass needle be deflected?

OpenStudy (nathan917):

a. d = 3 un = a + n − d u101 = 2 +100—3 u101 = 302 b. 152 = 2 + (n −1)(3) 150 = (n −1)(3), 50 = n −1, 152 = −1+ 3n n = 51